4. Suppose the wine industry is made up of many small identical firms. A represe tative firms long-run cost function is C(yi question 4 please.
Answer #1

(a)

Firm supply function is its MC curve.

MC = dC(y)/dy = - y + 3y2

So, firm supply curve: p = - y + 3y2

(b)

In long run equilibrium, p = MC = AC.

AC = C(y)/y = (99/2y) - (y/2) + y2

Equating with MC,

(99/2y) - (y/2) + y2 = - y + 3y2

99 - y2 + 2y3 = - 2y2 + 6y3

4y3 - y2 - 99 = 0

Solving this polynomial (using online solver),

y = 3 (other two roots are imaginary)

p = MC = - 3 + 3 x 3 x 3 = - 3 + 27 = 24

yD = 1,140 - 10 x 24 = 1,140 - 240 = 900

Number of firms = yD / y = 900 / 3 = 300

In long run equilibrium since p = AC, each firm earns zero profit.